Form the partial differential equation by eliminating the arbitrary constants a and b from z = xy + y((x+a) ^1/2) + b

1
Expert's answer
2022-03-31T14:01:57-0400

"z=xy+y\\sqrt{x+a}+b\\\\\\left\\{ \\begin{array}{c}\tz'_x=y+\\frac{y}{2\\sqrt{x+a}}\\\\\tz'_y=x+\\sqrt{x+a}\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\t\\sqrt{x+a}=z'_y-x\\\\\tz'_x=y+\\frac{y}{2\\left( z'_y-x \\right)}\\\\\\end{array} \\right. \\Rightarrow \\\\\\Rightarrow z'_x=y+\\frac{y}{2\\left( z'_y-x \\right)}"